OpenStudy (anonymous):

how do i figure out what's the hypotenuse of a right triangle when the base and height have a sq. root along side the number?

OpenStudy (psymon):

Do you have the specific numbers?

OpenStudy (anonymous):

the base and height is 5 sq. rooted

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

hypotenuse^2=base^2+height^2

OpenStudy (anonymous):

that's what i was thinking unkle but how would i go about doing it? im not too math smart, sorry.

OpenStudy (psymon):

Well, do you happen to know what exponent is the same as a square root?

OpenStudy (anonymous):

elaborate, please.

OpenStudy (psymon):

As in: \[\sqrt{5}=5^{\frac{ 1 }{ 2 }} \]

OpenStudy (anonymous):

so, i would set up the problem like so...\[5\frac{ 1 }{ 2 }^{2} + 5\frac{ 1 }{ 2 }^{2} = c ^{2} ??\]

OpenStudy (anonymous):

im trying to figure it out. im sorry.

OpenStudy (psymon):

A square root is the same as a 1/2 power. So when you need to square this, it looks like this: \[(5^{\frac{ 1 }{ 2 }})^{2}\] When you have an exponent that is being raised to another exponent, in this case an exponent of (1/2) that is being squared, you just multiply the exponents. So what is (1/2)*2?

OpenStudy (unklerhaukus):

\[(5^{1/2})^2=5^{2/2}=\]

OpenStudy (anonymous):

oooo.. i think im starting to get it. let me do this on paper, and ill post back my answer in a few seconds

OpenStudy (psymon):

Sure :3

OpenStudy (anonymous):

c would equal 10 ?

OpenStudy (unklerhaukus):

not quite

OpenStudy (anonymous):

i thought so

OpenStudy (anonymous):

didn't look right on paper.

OpenStudy (psymon):

You would have this: \[\sqrt{5^{\frac{ 1 }{ 2 }*2}+5^{\frac{ 1 }{ 2 }*2}}=\sqrt{5^{1}+5^{1}} \]

OpenStudy (anonymous):

would it then be 10 sq. ?

OpenStudy (psymon):

Correct :3

OpenStudy (anonymous):

YYYEEESSSSSSSS

OpenStudy (unklerhaukus):

yes, the hypo is √10 \(\checkmark\)

OpenStudy (anonymous):

thank you psymon and unkle ruck, what do i do now after i get help, do i hit the best response thingy?

OpenStudy (psymon):

Its not necessary xD It just gives one of us some medal thingy.

OpenStudy (unklerhaukus):

yes, hit best response if you want, then close the question, and then you will be able to open a new questioin

OpenStudy (anonymous):

aww man, alrighty. i still wanna thank you guys. im doing this quiz that ive been stuck on for a while because i was too afraid of failing again.

OpenStudy (psymon):

Well we always got people around to try and help :3

OpenStudy (unklerhaukus):

\[\color{orange}{\ddot\smile}\]

OpenStudy (anonymous):

awww man, the quiz told me it was wrong

OpenStudy (anonymous):

it* said it was 10 not 10 sq.

OpenStudy (unklerhaukus):

what?

OpenStudy (anonymous):

idk.

OpenStudy (psymon):

a^2 + b^2 = c^2 a = sqrt(5) b = sqrt(5) therefore \[\sqrt{5}^{2}+\sqrt{5}^{2}=c ^{2} \] \[10=c ^{2}->c=\sqrt{10}\]

OpenStudy (anonymous):

let me do this on paper. uno momento

OpenStudy (anonymous):

forgive me, laptop is slow, im typing out the equation

OpenStudy (unklerhaukus):

as you have worded the question above the hypotenuse is √10 as we have found. was the question in the quiz worded slightly differently ?

OpenStudy (anonymous):

\[5\sqrt{2}^{2} + 5\sqrt{2}^{2} = c ^{2}\]

OpenStudy (psymon):

Now thats 100% different.

OpenStudy (unklerhaukus):

oh right, do you mean ? \[(5\sqrt{2})^{2} + (5\sqrt{2})^{2} = c ^{2}\]

OpenStudy (psymon):

In that case it would be 10.

OpenStudy (anonymous):

i typed it before only with sq. root as 1/2 fraction, where was the error?

OpenStudy (psymon):

Its a 1/2 exponent. Just like squared is an exponent of 2, square root is an exponent of 1/2.

OpenStudy (anonymous):

alright, seeing that's the case, can you do the equation so i can see how it would be set up ?

OpenStudy (anonymous):

i learn from example

OpenStudy (psymon):

Well, the reason i wrote it as a 1/2 exponent was to help show you how you actually square it. But this is how I would go about it. Now remember, a square root is a 1/2exponent and if you square a square root, it just makes the square root go away. So this is how I'd write it: \[(5\sqrt{2})^{2}+(5\sqrt{2})^{2}=c ^{2}\]Now the key here is knowing I can rewrite the problem like this: \[5^{2}\sqrt{2}^{2}+5^{2}\sqrt{2}^{2}=c ^{2}\]You have to be aware that we are squaring everything. The 5 is squared as well as the sqrt(2). So this becomes: \[5^{2}2^{\frac{ 1 }{ 2 }*2}+5^{2}2^{\frac{ 1 }{ 2 }*2}\] \[(25)(2)^{1}+(25)(2)^{1}=c ^{2}\] \[50+50=c ^{2}->100=c ^{2}->\sqrt{100}=c->c=10\]

OpenStudy (anonymous):

the part where the 2 has the fraction and the exponent, can you explain to me how does that happen?

OpenStudy (psymon):

Right. Well remember \[\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\] The reason I write it like that is because then you can see that you have a 1/2 exponent being multiplied by a 2 exponent. Let's say I had this: \[(\sqrt{x})^{4} \] Well now we have a 4 power. It may not be obvious how to do that, but if we know that \[\sqrt{x}=x ^{\frac{ 1 }{ 2 }} \]then we can make more sense of this example. What we have is: \[x ^{\frac{ 1 }{ 2 }*4}=x ^{2} \]So the reason I put it as a 1/2 exponent is just to visually show you how we go about squaring, taking the 4th power, or doing anything to a square root.

OpenStudy (anonymous):

the symbol becomes the fraction?

OpenStudy (psymon):

Right. They are the same thing. \[\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\]A square root is just like having a 1/2 up there. Same thing. I just rewrote it as the fraction because it makes it easier to see what happens when you square it because it just becomes 1/2 * 2

OpenStudy (anonymous):

sorry ive been prolonging the convo, i just like multiple types clarification on things i dont grasp too quickly. now to be sure, can you give me an example problem?

OpenStudy (psymon):

You want to invent something maybe?

OpenStudy (anonymous):

literal or you mean something else?

OpenStudy (psymon):

Like invent a problem that you think might be good to have explained to you. So I can show you how I would handle all thwe square roots and powers.

OpenStudy (anonymous):

\[3\sqrt{5}^{2} + 7\sqrt{11}^{2} = c ^{2}\]

OpenStudy (psymon):

You mean for the 3 and the 7 to be squared, too, right?

OpenStudy (psymon):

I think you do, so Ill show ya.

OpenStudy (anonymous):

i hope my answer is the same as yours.

OpenStudy (psymon):

\[(3\sqrt{5})^{2}+(7\sqrt{11})^{2}\] So again, I want to make sure I square everything. That means I need to square the 3 and the sqrt(5) in the first problem. Square the 7 as well as the sqrt(11) in the 2nd one. So since I want to square everything, I can rewrite it like this: \[(3)^{2}(\sqrt{5})^{2}+(7)^{2}(\sqrt{11})^{2}=c ^{2}\]Now this is just so you can visually see what Im doing. Again, those square roots are just like having a 1/2 up there right next to the 2 power. So I'll rewrite it like that. \[(3)^{2}(5)^{\frac{ 1 }{ 2 }*2}+(7)^{2}(11)^{\frac{ 1 }{ 2 }*2}=c ^{2}\]Now having that 1/2 and that 2 up there, they get multiplied and just become 1. So now I have: \[(3)^{2}(5)+(7)^{2}(11)=c ^{2}\] 3 squared is 9 and 7 squared is 49, so I can fill those in: \[(9)(5)+(49)(11)=c ^{2}\] \[45+539=c ^{2}\] \[588=c ^{2}\] \[c=\sqrt{588}\] So that would be your answer. Have they shown you how to simplify radical expressions?

OpenStudy (anonymous):

it's 584 actually lol

OpenStudy (anonymous):

584 aq. rooted

OpenStudy (psymon):

Yeah, my bad, I didnt even look xD I just was typing as I was going xD

OpenStudy (anonymous):

lol, so i got it right this time

OpenStudy (psymon):

Awesome :D

OpenStudy (anonymous):

alright, thank you. im gonna try to finish my quiz now, thank you. i might come back for more mathematical wisdom soon.

OpenStudy (psymon):

Yeah, np. We'll be glad to help :3